OpenStudy (jaweria):

Please anyone help me with this question. Find Y prime: Y=sin^-1(2x^3).

OpenStudy (anonymous):

When you say y prime, do you mean the derivative?

OpenStudy (jaweria):

yup

OpenStudy (jaweria):

Please someone help me I m stuck and its due tomorrow

OpenStudy (goformit100):

Yo Yo, diffrentiate on both sides...

OpenStudy (jaweria):

These are kind of new problems to me if you dont mind can you teach me and help me solving them?

OpenStudy (anonymous):

the answer is:

OpenStudy (jaweria):

can we go step by step?

OpenStudy (anonymous):

Ya sure, I'll set it up.

OpenStudy (anonymous):

you want to go by steps sure will do

OpenStudy (anonymous):

at the end i will show u how u can work such equation in ur head, without paper and pen

OpenStudy (jaweria):

Thanks to both of you Sujay and Mathsmind :)

OpenStudy (anonymous):

Original Equation:\[y=\sin^2-1(2x-3)?\]

OpenStudy (anonymous):

with an x after the sin

OpenStudy (anonymous):

I want to make sure my original is correct

OpenStudy (anonymous):

\[y=\sin^{-1}(2x^3)\]

OpenStudy (anonymous):

please confirm is this what u r asking for?

OpenStudy (anonymous):

are u there?

OpenStudy (jaweria):

ok let me write down the equation here. Its \[y=\sin^{-1} (2x ^{3})\]

OpenStudy (anonymous):

i did write it

OpenStudy (jaweria):

yeah I m here sorry I was just writing that equation down

OpenStudy (jaweria):

yeah you are right Mathmind

OpenStudy (anonymous):

ok there are 2 main ways to solve this problem at ur level

OpenStudy (jaweria):

is there any easy way bcoz I have 4 more to do after this one so I need to understand the whole thing my professor method was too hard

OpenStudy (anonymous):

\[y = \arcsin(x)\]

OpenStudy (anonymous):

\[\sin(y)=x\]

OpenStudy (anonymous):

differentiate both sides

OpenStudy (anonymous):

Original:\[\sin^{-1} (2x^3)\] Use the chain rule: Derivative out the outside times the derivative of the inside. \[-\sin(2x^3)[\cos(2x^3)](6x^2) \] and there you have it, now simplify.

OpenStudy (anonymous):

\[cosy \frac{ dy }{ dx }=1\]

OpenStudy (anonymous):

\[\frac{ dy }{ dx }=\frac{ 1 }{ cosy }\]

OpenStudy (anonymous):

recall

OpenStudy (anonymous):

Apparently my answer is wrong, terribly sorry, I'll see where I went wrong :(.

OpenStudy (anonymous):

\[\sin^2y+\cos^2y=1\]

OpenStudy (anonymous):

Oh yes, my sine needs to be raised to a negative 2

OpenStudy (anonymous):

be patient both of u i will solve it in different way and u pick up the easiest method ok, all the best

OpenStudy (jaweria):

oh ok

OpenStudy (anonymous):

\[cosy=\pm \sqrt{1-\sin^2y}\]

OpenStudy (jaweria):

thank u so much Mathsmind and its ok Sujay

OpenStudy (anonymous):

\[\frac{ dy }{ dx } = \frac{ 1 }{ \sqrt{1-\sin^2y} }\]

OpenStudy (anonymous):

now remember our rearrangement\[\sin(y)=x\] so substitute that in the above equation

OpenStudy (anonymous):

\[\frac{ dy }{ dx } = \frac{ 1 }{ \sqrt{1-x^2} }\]

OpenStudy (anonymous):

now this is the general derivative of arcsin or sin inverse ok

OpenStudy (anonymous):

now if you follow the same steps for\[y = \sin^{-1}(2x^3)\]

OpenStudy (anonymous):

\[\frac{ dy }{ dx } = \frac{ 6x^2 }{ \sqrt{1-4x^6} }\]

OpenStudy (anonymous):

and that's is your final answer achieved using the chain rule

OpenStudy (anonymous):

let me prove that for you

OpenStudy (jaweria):

ok plzz if you can do that

OpenStudy (anonymous):

step One:\[2x^3=siny\]

OpenStudy (anonymous):

step two: differentiate both sides

OpenStudy (anonymous):

\[6x^2=\frac{ dy }{ dx }cosy\]

OpenStudy (anonymous):

step three : rearrange the differential equation

OpenStudy (anonymous):

\[\frac{ dy }{ dx}=\frac{ 6x^2 }{ cosy }\]

OpenStudy (anonymous):

step four: recall that \[cosy=\pm \sqrt{1-\sin^2y}\]

OpenStudy (anonymous):

step five: substitute cosy in the differential equation \[\frac{ dy }{ dx}=\frac{ 6x^2 }{ \sqrt{1-\sin^2y} }\]

OpenStudy (jaweria):

wow I am liking this method better than before

OpenStudy (anonymous):

step six: recall\[siny=2x^3s\]

OpenStudy (anonymous):

Step seven: substitute siny in the last deferential

OpenStudy (jaweria):

what number is that under the power of 3?

OpenStudy (anonymous):

that is a mistake an s

OpenStudy (anonymous):

sorry i will do it again

OpenStudy (jaweria):

oh ok

OpenStudy (anonymous):

back to step six ok

OpenStudy (anonymous):

\[siny=2x^3\]

OpenStudy (jaweria):

ok

OpenStudy (anonymous):

step seven : substitute step 6 in step 5

OpenStudy (anonymous):

\[\frac{ dy }{ dx}=\frac{ 6x^2 }{ \sqrt{1-(2x^3)^2} }\]

OpenStudy (anonymous):

finally step eight : simplify

OpenStudy (anonymous):

\[\frac{ dy }{ dx}=\frac{ 6x^2 }{ \sqrt{1-4x^6} }\]

OpenStudy (anonymous):

and this concludes our proof good luck

OpenStudy (anonymous):

are you there? am sorry my computer is slow today

OpenStudy (jaweria):

oh wow you explained really good thanks :)

OpenStudy (anonymous):

so i deserve a medal then hehehe

OpenStudy (jaweria):

Now I have to work on 4 more of these huh lets see how am I going to do it.

OpenStudy (jaweria):

yup you did :)

OpenStudy (anonymous):

ok if you do those 4 by yourself i will give you a medal, does that sound fair?

OpenStudy (jaweria):

yup it does :)

OpenStudy (anonymous):

you can use the triangle method as well to solve this

OpenStudy (jaweria):

if you dont mind can I just step by step show you how to do that?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

how am i gonna give u a medal then ! hehehe

OpenStudy (jaweria):

thanks

OpenStudy (jaweria):

oh yeah sorry i forgot hehe

OpenStudy (anonymous):

ur welcome!

OpenStudy (anonymous):

u better not make one mistake

OpenStudy (jaweria):

ohhh hmmm I am not sure about that but I m doing it

OpenStudy (anonymous):

all the best

OpenStudy (anonymous):

there is another method using Taylor's series

OpenStudy (jaweria):

my professor also gave us the formula like for that equation the formula for \[\sin^{-1} is 1/\sqrt{1-u ^{2}}timesdu\]

OpenStudy (anonymous):

yes that is the method of substitution which i see it as a lazy method

OpenStudy (jaweria):

oh ok

OpenStudy (anonymous):

but if u are used to that method then go for it, as soon as u get the right answer who cares, you know!

OpenStudy (jaweria):

yeah

OpenStudy (jaweria):

i m confuse

OpenStudy (jaweria):

the answer that I m getting is weird

OpenStudy (jaweria):

are u there?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

post the question

OpenStudy (anonymous):

r u there?